# State-of-the-Art Dysarthric Speech Recognition with MetaICL for on-the-fly Personalization

### v1

@@ Abstract
Personalizing Automatic Speech Recognition (ASR) for dysarthric speech is crucial but challenging due to training and storing of individual user adapters. We propose a hybrid meta-training method for a single model, excelling in zero-shot and few-shot on-the-fly personalization via in-context learning (ICL). Measuring Word Error Rate (WER) on state-of-the-art subsets, the model achieves 13.9% WER on Euphonia which surpasses speaker-independent baselines (17.5% WER) and rivals user-specific personalized models. On SAP Test 1, its 5.3% WER significantly bests the 8% from even personalized adapters. We also demonstrate the importance of example curation, where an oracle text-similarity method shows 5 curated examples can achieve performance similar to 19 randomly selected ones, highlighting a key area for future efficiency gains. Finally, we conduct data ablations to measure the data efficiency of this approach. This work presents a practical, scalable, and personalized solution.
@@ 1.1 Motivation
The acoustic variability and reduced intelligibility of dysarthric speech presents significant challenges for Automatic Speech Recognition (ASR) systems [1]. While personalization is a critical step to bridge this performance gap, conventional approaches are often burdensome. Methods like user-speaker fine-tuning or deploying parameter-efficient adapters [2], while effective, require training and storing separate models for each user, posing significant complexities for large-scale deployment and maintenance. In-context learning (ICL) has emerged as a paradigm shift, enabling personalization [3] purely at inference time by providing a few user-specific examples. Specifically for ASR, we would include many audio-text examples (“shots”) to support the query example. This eliminates the need for per-user training, motivating the search for a truly universal model that can serve and personalize to any speaker on-the-fly.
@@ 1.2 Related Work
As discussed, Dysarthric ASR (DSR) research has explored fine-tuning and parameter-efficient methods like LoRA for individual speaker adaptation, with studies demonstrating personalization is possible even with small datasets (e.g.,  250 phrases per user in [4]). Recent advancements include personalized RNN-T [5] approaches outperforming speaker-independent models by 35% [6] on Euphonia [7], and AdaLoRA [8] improving over fine-tuning by 24% [9] on the Speech Accessibility Project dataset [10]. While these methods have pushed the state-of-the-art (SOTA), reducing the memory overheads in adapters is an active area of research [11].
The emergence of large multi-modal models offers a new avenue: inference-time personalization through meta-learning or “MetaICL” [12], which uses contextual examples for on-the-fly personalization without weight updates. This work investigates MetaICL’s applicability and scalability for dysarthric ASR personalization across diverse datasets, aiming for state-of-the-art performance. It also investigates if a single model performs well with and without few-shot examples, analyzes MetaICL’s influence on learning dynamics, and examines the role of example curation [13] for efficiency in ASR, similar to its effect on machine translation [14].
@@ 1.3 Our Contributions
This paper introduces a hybrid meta-training strategy to create a single, universal model that achieves SOTA performance and is highly practical to deploy. Our key contributions are:
@@
    A Novel Hybrid Training Strategy: We propose and validate a mixed training setup using both 0-shot (query only) and 10-shot (support + query) examples, resulting in a model that excels in all evaluation scenarios.
@@
    Scalable Across Diverse Datasets: First to demonstrate the effectiveness of this approach on both – Euphonia and Speech Accessibility Project (SAP). We establish new SOTA benchmarks and prove the scalability of this technique to different speech impairments.
@@
    Effective Example Curation: We explore different few-shot example curation methods via text-based embeddings and show how it can significantly reduce the number of user examples needed at inference.
@@
    Data Ablation Study: We explore the learning dynamics of our approach, with majority of performance gains coming from a fraction of data.
@@ 2 Methodology
Fig. 1: Method Overview. Left: Multi-modal LLM architecture. Middle: Example prompt construction. Right: Mixed (0-shot, 10-shot) MetaICL training & evaluation with example curation.
@@
Our work utilizes the Gemini 2.5 Flash model [15], a multimodal LLM with advanced audio understanding and instruction-following capabilities. Its strong baseline instruction following performance aids the model to have strong capacity to learn with meta-learning fine-tuning. We only train the LLM layers while using the frozen audio tokenizer. The base model has not been trained on dysarthric speech.
@@ 2.2 Meta-Learning for In-Context Personalization
We frame the personalization task as a meta-learning problem. The goal is to train a single model that, at inference time, can learn to recognize a specific user’s speech patterns from a small set of support examples provided in the prompt. An inference-time prompt consists of:
@@
The model’s task is to predict the transcript $t_{q}$ for query $a_{q}$, conditioned on the support set. The training process optimizes the model’s ability to perform in-context adaptation.
@@ 2.3 Datasets and Evaluation
We conduct experiments on two of the largest publicly available dysarthric speech datasets, ensuring our splits maintain speaker and phrase separation.
  • Google’s Euphonia Project: A large-scale research initiative containing audio samples from more than 2,000 speakers with diverse etiologies (e.g., ALS, Parkinson’s, vocal cord paralysis). For our experiments we have limited to using approx 300k utterances (about a third) in the training dataset. The test-set contains 5,684 utterances from over 350 speakers, reviewed by speech-language pathologists. For fair SOTA comparisons, we use a subset of 1,740 utterances.
  • Speech Accessibility Project (SAP): A collaboration led by the University of Illinois Urbana-Champaign and major tech companies to create a shared dataset for research. The training set contains 240k utterances, and the two distinct test sets (Test1 and Test2) have approximately 35k utterances each. For SOTA evaluation conditions, we have a subset of roughly 17k utterances from Test1 to match the dataset used for Interspeech 2025 [16].
Table 1: WER (%) comparison of different training strategies. Best results for each n-shot condition and dataset are in bold.
Model (Training Data) Euphonia SAP-Test1 SAP-Test2
@@
Table 2: SOTA Comparison on Euphonia and SAP published benchmark subsets. All results are WER (%). Lower is better.
Evaluation Set Split Our Model (N-shot ICL) Non-personalized SOTA Personalized SOTA
@@
SAP Test1 (Subset) Overall 7.5 (0-shot); 5.3 (10-shot) 10.6 (FFT) [8] 8 (AdaLORA) [8]
SAP Test2 (Subset) Overall 10.1 (10-shot) 10.0 (Whisper Self-Training) [17] -
Datasets vary in dysarthria etiology and severity distributions. Performance is reported as Word Error Rate (WER %).
@@ 2.4 Training Strategies
A key contribution of this work is our investigation into the optimal training strategy with these four approaches:
@@
    0-Shot Trained (SFT): This approach represents the baseline traditional way of training a LLM-based ASR model without explicitly teaching it to use context examples. We train on a large corpus of speech (200k examples). Each example consists of a single audio-text pair, $(a_{q},t_{q})$.
@@
    Mixed-Objective (0+10 Shot) Trained: This is our proposed solution. We train the model on a mixture of 10-shot and 0-shot (upto 40k each) examples. We hypothesize that this model learns both to perform high-quality zero-shot transcription and to effectively utilize contextual examples.
2.5 Text-based Enrollment Curation
Our default method uses random utterances from the speaker to be used as support examples. Inspired by machine translation work [14, 18], we explore if curating ICL examples using text-based methods improves the efficiency of meta-learned ASR models. First, we obtain sentence embeddings of transcripts, using the publicly available Universal Sentence Encoder model [19]. Based on these sentence embeddings, we compare the below sampling methods:
@@ 2.5 Text-based Enrollment Curation
    Random: Sample random utterances from the same user as support examples (default).
@@
    Diversity (Centroid): Builds clusters based on transcript sentence embeddings. For 5-shot, uses 5 cluster centroids as the representative ICL examples for the user to provide diverse phonetic and lexical range.
@@
    Sorted by Uncertainty: Uses active learning concepts to select challenging utterances as ICL examples. Performs 0-shot evaluation, then selects utterances with the highest WER.
  4. 4.
    Sorted by Similarity: Finds utterances most textually similar to the current query by computing cosine similarity between sentence embeddings, then selects those with the highest similarity. We include this method as an oracle experiment. Since it requires the query’s transcript, it is not viable for real-world deployment. Its purpose is to demonstrate the potential upper bound of curation, motivating future work into practical alternatives like acoustic embedding similarity.
Table 3: Comparison of 5-shot curation strategies against random curated 19-shot on Euphonia. Metric: WER (%).
Curation Strategies (5-shot) WER (%)
Random 11.3
Diversity (Cluster Centroid) 11.0
Sorted (desc.) by Uncertainty 11.4
Sorted (desc.) by Similarity 9.9
19-shot Random Baseline 9.5
@@ 3.1 Training Strategy Comparison
Our primary experiments, detailed in Table 1, evaluate the efficacy of different training strategies. The results demonstrate the superiority of a mixed-objective approach.
@@ 3.1.1 Comparing Single-Objective Training on Euphonia
When training exclusively on Euphonia, we observe a clear trade-off with single-objective models. The 0-shot model is trained on 200k examples with 1 utterance each, compared to 20k examples with 11 utterances each for 10-shot. The 0-shot model achieves a 0-shot relative reduction of over 52% from baseline in WER but is unreliable at leveraging ICL examples, with performance degrading from 5-shot to 10-shot. Conversely, the 10-shot only model excels at personalization, reaching an impressive 9.9% WER at 19 shots, but it’s 0-shot WER of 24.4% is significantly weaker. This highlights that single objective models are not optimal for both scenarios.
@@ 3.1.2 Mixed-Objective and Multi-Dataset Synergy
The next model, Euphonia based mixed-objective model, extends the 10-shot model with additional 20k examples for 0-shot . Successfully resolving the trade-off, it achieves a strong 0-shot WER of 17.1%, nearly matching the 200k 0-shot only model, while having excellent and consistent few-shot performance (9.8% at 19 shots). The model generalizes impressively, reducing the SAP-Test1 0-shot WER by 36.6% without training on it. Also, the consistent N-shot performance shows we don’t need to train on varying count of support examples.
The advantages are further amplified when incorporating SAP data into the training mix in our final model, achieving the best results across all conditions. On Euphonia, the 0-shot WER improves to 16.3% and the 19-shot WER to 9.5%. Similar improvements on SAP, with the SAP-Test1 0-shot WER plummeting to 11.5% and the 10-shot WER to 8.8%. This demonstrates that a mixed-objective, mixed-dataset approach yields the most robust and personalized DSR system.
@@ 3.2 State-of-the-Art Comparison
As shown in Table 2, our best model significantly advances the SOTA on standardized subsets. On Euphonia, we achieve dramatic WER reductions across all severity levels compared to the USM [20] SI-ASR speaker-independent baseline [6], with WER for moderate dysarthria dropping from 19.6% to just 11.2%. On SAP Test1, our model is a clear leader, achieving a 5.3% WER with 10 shots, outperforming even personalized SOTA methods of AdaLORA (8%) [8]. While a fully personalized RNN-T model still achieves a lower WER on the Euphonia subset (11.3%) [6], our model (13.9%) surpasses speaker-independent baselines (17.5% WER). Reinforcing that our approach offers competitive performance from a single, universal model with zero user-specific training.
3.3 Enrollment Curation
Given the impracticality of requiring numerous enrollment examples, we explored potential value of curation strategies to demonstrate their importance on the best model from Table 1. As shown in Table 3, we tested several methods, including an ’oracle’ benchmark based on textual similarity to the query (as described in Section 2.5). While this specific method is not feasible in deployment, its 5-shot performance (9.9% WER) nearly matched the 19-shot random baseline (9.5% WER). This result’s primary importance is its demonstration of curation’s potential efficiency—showing 5 well-chosen examples could replace 19 random ones—and establishing it as a critical area for improving the on-the-fly experience.
Interestingly, strategies based on diversity and uncertainty did not yield strong improvements. This may suggest that for a given speaker, the semantic content within their enrollment utterances lacks sufficient diversity for a clustering approach to be beneficial.
3.4 Data Ablation Study
To understand the learning dynamics of our mixed-objective approach, we conducted a data ablation study, with total data of 40k 10-shot data and 10k 0-shot. Our study (Table 4) reveals a highly front-loaded learning curve. The initial 2% of training data providing over 70% of the total possible WER reduction. This rapid learning stems from the data-rich meta-learning format; while the loss is only calculated on the single query utterance, the model learns crucial speaker traits from the 10 support examples during the forward pass.
We also observe diminishing returns, with gains slowing after the 40% data mark. This suggests that near-SOTA performance is achievable with less than half the training set, highlighting that our model learns from MetaICL quickly.

### v2

# The Universal Personalizer: Few-Shot Dysarthric Speech Recognition via Meta-Learning

@@ Abstract
Personalizing dysarthric ASR is hindered by demanding enrollment collection and per-user training. We propose a hybrid meta-training method for a single model, enabling zero-shot and few-shot on-the-fly personalization via in-context learning (ICL). On Euphonia, it achieves 13.9% Word Error Rate (WER), surpassing speaker-independent baselines (17.5%). On SAP Test-1, our 5.3% WER outperforms the challenge-winning team (5.97%). On Test-2, our 9.49% trails only the winner (8.11%) but without relying on techniques like offline model-merging or custom audio chunking. Curation yields a 40% WER reduction using random same-speaker examples, validating active personalization. While static text curation fails to beat this baseline, oracle similarity reveals substantial headroom, highlighting dynamic acoustic retrieval as the next frontier. Data ablations confirm rapid low-resource speaker adaptation, establishing the model as a practical personalized solution.
@@ 1.1 Motivation
Acoustic variability in dysarthric speech  [21] creates a ‘utility gap’ that standard ASR systems [2] cannot bridge. Traditional personalization carries a prohibitive human cost: users with limited vocal stamina must often record hundreds of phrases to train dedicated models or adapters [17], creating a barrier for those with degenerative conditions who need the technology most. Aligning with the Interspeech 2026 theme of ‘Speaking Together,’ we propose a ‘Universal Personalizer’—a system that understands a unique voice in seconds, purely through context. By re-framing personalization as an inference-time task via In-Context Learning (ICL) [13], we provide a few audio-text ‘shots’ to guide the model on-the-fly. This eliminates per-user training and gradient updates, enabling a truly universal model that provides immediate, high-quality ASR without the friction of data-heavy enrollment.
@@ 1.2 Related Work
Dysarthric ASR (DSR) research [11] has explored fine-tuning and parameter-efficient methods like LoRA for individual speaker adaptation, with studies demonstrating personalization is possible even with small datasets (e.g.,  250 phrases per user in [16]). Recent advancements include personalized RNN-T [7] approaches outperforming speaker-independent models by 35% [14] on Euphonia [9], and AdaLoRA [18] improving over fine-tuning by 24% [22] on the Speech Accessibility Project dataset [8]. Recently, the Interspeech 2025 SAP Challenge established new benchmarks. Top-performing systems achieved these results through strategies like offline weight-merging and custom audio segmentation. While these optimization techniques and adapters effectively push performance boundaries, large multi-modal models offer a distinct alternative paradigm: inference-time adaptation through “MetaICL” [10] or meta-learning.
Our approach uses meta-learning to use contextual examples for on-the-fly personalization [13] without weight updates. This work investigates its applicability and scalability for dysarthric ASR personalization across diverse datasets, aiming for state-of-the-art performance. It also explores if a single model performs well with and without few-shot examples, analyzes MetaICL’s influence on learning dynamics, and examines the role of example curation [19] for efficiency in ASR, similar to its effect on machine translation [1].
@@ 1.3 Our Contributions
The core contributions discussed in the paper are majorly:
@@
    Hybrid Training Strategy: We propose and validate a mixed training setup using both 0-shot (query only) and 10-shot (support + query) examples, excelling in all evaluation setups.
@@
    Effectiveness on Diverse Datasets: We test this on the biggest dysarthric sets – Euphonia and Speech Accessibility Project (SAP). We set new non-personalized SOTA benchmarks and narrow the performance gap with per-user models.
@@
    Effective Example Curation: Demonstrate how our trained model leverages same-speaker support examples to do on-the-fly personalization. We also explore different example curation methods via text-based embeddings, while establishing a theoretical upper bound on curation.
@@
    Data Ablation: We analyze learning dynamics, specifically the rates of speaker and domain adaptation.
@@ 2 Methodology
Fig. 1: Overview. Left to Right: Architecture; Example construction; Mixed MetaICL training & evaluation with example curation.
@@
Our work utilizes the Gemini 2.5 Flash model [4], a multi-modal LLM with advanced audio understanding and instruction-following capabilities. Its strong baseline instruction following performance aids the model to have strong capacity to learn with meta-learning fine-tuning. We only train the LLM layers while using the frozen audio tokenizer. The base model has not been trained on dysarthric speech. While we utilize Gemini 2.5 Flash for its incredible audio understanding capabilities, our MetaICL framework is model-agnostic and transferable to other multi-modal LLMs (e.g., Qwen-ASR[12], Llama-Omni 2 [5]) for similar on-the-fly personalization capabilities. However, our goal here was to showcase applicability and achieve best performance.
@@ 2.2 Meta-Learning for In-Context Personalization
We frame personalization as meta-learning, training a single model to recognize user-specific speech patterns via a small set of support examples provided in the prompt. Our prompt consists of:
@@
The model’s task is to predict the transcript $t_{q}$ for query $a_{q}$, conditioned on the support set. The loss is calculated only on the query utterance. The training process optimizes the model’s ability to perform in-context adaptation.
@@ 2.3 Datasets and Evaluation
We use two of the largest available dysarthric speech datasets.
  • Google’s Euphonia Project: A large research initiative containing audio samples from more than 2,000 speakers with diverse etiologies (e.g., ALS, Parkinson’s, vocal cord paralysis). We limited training to approximately 300k utterances (almost 30% of the total). The test-set contains 5,684 utterances from over 199 speakers, reviewed by speech-language pathologists. For fair SOTA comparisons, we use a subset of 1,740 utterances.
  • Speech Accessibility Project (SAP): A collaboration led by the University of Illinois Urbana-Champaign and major tech companies to create a shared dataset for research. The training set contains 240k utterances, and the two distinct test sets (Test-1 and Test-2) have approximately 35k utterances each. For SOTA evaluation, we utilize the 17k-utterance Test-1 subset established by the Interspeech 2025 Speech Accessibility Project (SAP) Challenge [24]. Similar to Euphonia, we enforce strict speaker separation between splits.
Table 1: WER (%) comparison of different training strategies. We use complete test data available for Euphonia and SAP here.
Model (Training Data) Euphonia SAP-Test-1 SAP-Test-2
@@
Table 2: Comparisons on published versions of test-sets, Euphonia and SAP, w.r.t speaker independent and dependent benchmarks.
Evaluation Set Split Our Model (N-shot ICL) Non-personalized SOTA Per-user SOTA
@@
SAP Test-1 (Subset) Overall 7.5 (0-shot); 5.3 (10-shot) 5.9 (Parakeet-tdt [24]) 10.6 (FFT) [18] 8 (AdaLoRA) [18]
SAP Test-2 (Subset) Overall 9.5 (10-shot) 8.1(Parakeet-tdt [24]) 10.0 (self-training) [20] -
We ensure strictly disjoint speakers between training and test sets to prevent information leakage. Diversity of dysarthria etiology and severity distributions is discussed in ([9], [15], [24]). Performance is reported as Word Error Rate (WER %).
@@ 2.4 Training Strategies
We investigate four training strategies:
@@
    0-Shot Trained (SFT): This represents standard LLM-based ASR training without explicit context utilization. We train on a corpus of 200k examples of single audio-text pairs, $(a_{q},t_{q})$.
@@
    Mixed-Objective (0+10 Shot) Trained: Our proposed solution, where we train the model on a mixture of 10-shot and 0-shot examples. We hypothesize that this model learns both to perform high-quality zero-shot transcription and to effectively utilize contextual examples.
2.5 Optimization of In-Context Example Selection
Our default approach utilizes random utterances from the target speaker as support examples. Inspired by machine translation work [1, 6], we explore if curating ICL examples influences ICL performance. We investigate three categories of curation strategies to optimize this selection:
Table 3: Curation strategies comparison on Euphonia. Only “Random Different Speaker“ doesn’t use same-speaker examples.
Curation Strategies (5-shot) Type WER (%)
Random (Different Speaker) Baseline 18.8
Random (Same-Speaker) Our Default 11.3
Diversity (Cluster Centroid) Offline/Static 11.0
Most Uncertain Offline/Static 11.4
Most Similar Oracle (Dynamic indexing) 9.9
19-shot Random (same speaker) Upper Baseline 9.5
@@ 2.5 Optimization of In-Context Example Selection
    Random: We compare support sets comprised of (a) Random (different speaker) utterances against (b) Random (same speaker) utterances; to isolate the impact of personalization on speaker-specific acoustics.
@@
    Computed Static Representative Sets: We propose ”offline” selection methods that pre-calculate a single, optimal set of $N$ examples per user to serve as a static representation of their speech. We do this using text-based methods on a set of utterances per user (enrollment set). First, we obtain sentence embeddings of transcripts, using the publicly available Universal Sentence Encoder model [3].
    1. (a)
      Diversity (Centroid): We cluster the user’s sentence embeddings and select cluster centroids to maximize phonetic and lexical coverage within the $N$-shot support set.
    2. (b)
      Sorted by Uncertainty: We perform a 0-shot pass on the user’s enrollment data and select utterances with the highest WER. The rationale is that providing the model with its own ”weak spots” (hard negatives) may better calibrate it to the user’s specific dysarthric patterns.
@@
    Oracle with text-similarity (Dynamic): While not deployable (as it requires the ground truth), this method establishes the potential theoretical lower bound of WER achievable through optimal example retrieval. For each query, we select support examples with the highest cosine similarity between their transcript embeddings and the query transcript.
@@ 3.1 Training Strategy Comparison
Our primary experiments, detailed in Table 1, evaluate the efficacy of different training strategies.
@@ 3.1.1 Comparing Single-Objective Training on Euphonia
When training exclusively on Euphonia, we observe a clear trade-off with single-objective models. The 0-shot model is trained on 200k examples with 1 utterance each, compared to 20k examples with 11 utterances each for 10-shot. The 0-shot model reduces baseline WER by 52% but fails to leverage ICL examples, with performance degrading from 5-shot to 10-shot. Conversely, the 10-shot only model excels at personalization, reaching an impressive 9.9% WER at 19 shots, but it’s 0-shot WER of 24.4% is significantly weaker. This highlights that single objective models are not optimal for both scenarios.
@@ 3.1.2 Mixed-Objective and Multi-Dataset Synergy
The next model, Euphonia based mixed-objective model, extends the 10-shot model with additional 20k examples for 0-shot . Successfully resolving the trade-off, it achieves a strong 0-shot WER of 17.1%, nearly matching the 200k 0-shot only model, while having excellent and consistent few-shot performance (9.8% at 19 shots). The model generalizes impressively, reducing the SAP-Test-1 0-shot WER by 36.6% without training on it. Also, the consistent N-shot performance shows we don’t need to train on varying count of support examples.
The advantages are further amplified when incorporating SAP data into the training mix in our final model, achieving the best results across all conditions. On Euphonia, the 0-shot WER improves to 16.3% and the 19-shot WER to 9.5%. Similar improvements on SAP, with the SAP-Test-1 0-shot WER plummeting to 11.5% and the 10-shot WER to 8.8%. This demonstrates that a mixed-objective, mixed-dataset approach yields the most robust and personalized DSR system.
@@ 3.2 State-of-the-Art Comparison
As shown in Table 2, our best model significantly advances the SOTA on standardized subsets. On Euphonia, we achieve dramatic WER reductions across all severity levels compared to the USM [23] SI-ASR speaker-independent baseline [14], with WER for moderate dysarthria dropping from 19.6% to just 11.2%. While a fully personalized RNN-T model still achieves a lower WER on the Euphonia (11.3%) [14], our model (13.9%) surpasses speaker-independent baselines (17.5% WER). On SAP Test-1, our model establishes a new state-of-the-art, achieving a 5.3% WER (10-shot). This outperforms both previous personalized methods like AdaLoRA (8.0%) [18] and the Interspeech 2025 SAP Challenge winning system (5.97% [24]). On SAP Test-2, our model achieves 9.5% WER. While this trails the absolute challenge winner (8.11% [24]), it comfortably surpasses the second-best submission (10.0% [20]). Crucially, our approach achieves this top-tier performance on both sets using a single, unified architecture, completely avoiding techniques like complex model-merging and custom audio chunking utilized by the challenge winners.
3.3 Efficacy of Enrollment Curation
We conduct the experiments on the best model from Table 1 and discuss the results shown in Table 3 in-depth below:
3.3.1 Validating On-the-Fly Personalization
The personalization mechanism is isolated by comparing random support utterances from same-speaker versus different-speaker sets.
Providing 5 mismatched support examples yields 18.8% WER, a degradation compared to the 16.3% 0-shot baseline from Table 1. This “interference effect” confirms the model actively maps support acoustics to the query, failing when identities conflict. Conversely, same-speaker examples drop WER to 11.3%. This relative 40% drop validates that the model utilizes in-context audio for precise speaker adaptation rather than generic domain robustness.
3.3.2 Static Curation Strategies
We evaluated if ”Computed Static Sets” (Diversity and Uncertainty) could outperform random selection (Table 3). We find that random same-speaker selection is a robust baseline (11.3%) which static curation fails to surpass. This indicates the primary driver of personalization here is acoustic consistency rather than the semantic diversity of the support set.
3.3.3 The Efficiency Headroom
Finally, our oracle text-similarity experiment demonstrates the theoretical upper bound of our approach. Semantically similar 5-shot examples reduces WER to 9.9%, close to the performance of a 19-shot random baseline (9.5%) with only a quarter of the few-shot data. This result highlights a significant ”efficiency headroom.” It indicates that while text-based static curation methods have saturated, dynamic retrieval strategies hold promise. This motivates future work into audio-based similarity indexing, which could approximate this oracle performance in a deployable, text-free manner.
3.4 Data Ablation: Domain vs. Speaker adaptation comparison
We conducted a data ablation study (Table 4), with 50k instances (40k 10-shot and 10k 0-shot). The loss is calculated only on the query utterance. The 4:1 ratio prioritizes the acquisition of the in-context learning mechanism while preserving sufficient zero-shot supervision and being optimal in training steps needed overall. We see a distinct decoupling between the speaker and domain adaptation.
  • Rapid Mechanism Acquisition (Speaker Adaptation): With just 2% or 1000 instances (800 10-shot & 200 0-shot), the gap between 0-shot (22.6%) and 10-shot (14.8%) is fully established. This suggests that the logic of attending to user-specific context—is highly data-efficient and established early in training.
  • Domain Adaptation comes next: In contrast, the 0-shot baseline improves till 40% of data at 17.8% (or 16k 10-shot & 4k 0-shot examples) and then plateaus. This indicates that domain adaptation needed more training rounds but less than 50% of our initial data is needed to have a competitive model.
This decoupling highlights that near-SOTA personalization capabilities can be achieved with a fraction of the data, provided the model has sufficient exposure to learn the ICL interaction pattern.
3.5 Inference Limitations vs. Operational Efficiency
We acknowledge that multi-modal LLM backbones require higher inference costs than lightweight RNN-T or LoRA-based models. However, user studies indicate that moderate-to-severe dysarthric speakers currently lack any functional ASR solutions. By achieving 11.2% WER for moderate severity, we prioritize crossing this critical ”usability threshold” over latency minimization, as the gain in fundamental utility outweighs the inference cost. Additionally, this approach optimizes operational efficiency by eliminating per-user training infrastructure, aligning with the foundation model paradigm: a single, universal model capable of instant, stateless personalization without the complexity of gradient-based updates.